OpenStudy (anonymous):

The students at Midtown Middle school sold flowers as a fundraiser in September and October. In October, they charged $1.50 for each flower. The October price was a 20% increase of the September price. Part A What was the price of the flowers in September? Part B The seventh-grade class earned 40% of the selling price of each flower. In September, they sold 900 flowers. In October, they sold 700 flowers. Select a choice from each drop-down menu to make a true statement. The seventh-grade class earned more money in......OCTOBER/SEPTEMBER.....by.......$30/$75/ $190.

OpenStudy (anonymous):

Part A: Let's say september price is x That means: x+0.2x=1.5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I don't know how to explain. Can I tell you the answer though?

OpenStudy (anonymous):

um sure i'll just figure it out afterwards

OpenStudy (anonymous):

It's 1.25. Checking it over, 20% of 1.25 is 0.25, and 1.25+0.25=1.50, so it's correct.

OpenStudy (anonymous):

ok i kinda knew that part but what in stuck on is part

OpenStudy (anonymous):

b

OpenStudy (anonymous):

I don't really get what they mean by they earned 40% of the selling price of each flower.

OpenStudy (anonymous):

But, I think it means that they sold the flowers 40% more than the price they bought it.

OpenStudy (anonymous):

Let me think it over some more.

OpenStudy (anonymous):

i thought it meant from the total they earned 40% of it

OpenStudy (anonymous):

I think you are correct.

OpenStudy (anonymous):

So then, multiply the cost of the flowers by the number of flowers they sold and by 0.4 for each month and compare.

OpenStudy (anonymous):

Can you write it in an expression plz? lol

OpenStudy (anonymous):

ok. 0.4*(1.5*700) for october o.4*(1.25*900) for september

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

ur welcome.

OpenStudy (anonymous):

420 for october

OpenStudy (anonymous):

450 for september?

OpenStudy (anonymous):

@jf04

OpenStudy (anonymous):

@Baby_Bear69 @DullJackel09 @jasmineETTA @HelloKitty17

OpenStudy (anonymous):

@WhateverYouSay

OpenStudy (anonymous):

@robtobey @Robert136

OpenStudy (anonymous):

@jay231

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

@acero1984

OpenStudy (anonymous):

@Elsa213 @RedNeckOutLaw @foreverthebeast @GIL.ojei @enchanted_bubbles

OpenStudy (anonymous):

jfo4s answer was all wrong...

OpenStudy (anonymous):

Oh. Thats nice lol.

OpenStudy (anonymous):

very lol

OpenStudy (anonymous):

Give me 1 min.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

@AdamLevine1986

OpenStudy (anonymous):

Okay... Oct -- 1.50 per flower Sept -- 20% of 1.50 = 0.2(1.5) = 0.30 : 1.50 - 1.30 = 1.25 per flower Part B : Sept : sold 900 flowers at 1.20 .....900 x 1.20 = 1080 40% of 1080 = (0.4)(1080) = 432 Oct : sold 700 flowers at 1.50...700 x 1.50 = 1050 40% of 1050 = (0.4)(1050) = 420 earned more money in Sept. For more details you can check out these websites:] http://openstudy.com/study#/updates/548a08f0e4b0f2fbdc2184f1 http://openstudy.com/study#/updates/53f508e9e4b0f0e909dfec0c Good luck!

OpenStudy (anonymous):

wait i checked those, they were wrong too

OpenStudy (anonymous):

Oh hmmmm

OpenStudy (anonymous):

Thats weird...

OpenStudy (anonymous):

ik..

OpenStudy (anonymous):

:|

OpenStudy (anonymous):

this question is stupid

OpenStudy (anonymous):

Have you tried 1.20 ?

OpenStudy (anonymous):

i tried that and 1.25

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

Give me 1 min!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What about 1.30?

OpenStudy (anonymous):

Hmmm let me check...

OpenStudy (anonymous):

Are you in FLVS?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Algebra 1?

OpenStudy (anonymous):

i honestly dont even know

OpenStudy (anonymous):

LOL.

OpenStudy (anonymous):

Well if you tell me and the lesson number i could help!

OpenStudy (anonymous):

4.08

OpenStudy (anonymous):

ok give me 1 min.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I can't seem to find it i am sorry :-|

OpenStudy (anonymous):

Good luck!

OpenStudy (anonymous):

bye i guess

OpenStudy (anonymous):

anyone??

OpenStudy (acero1984):

I can't find it

OpenStudy (anonymous):

ik i cant either

OpenStudy (acero1984):

:'(

OpenStudy (anonymous):

...you should give me a medal for trying lol

OpenStudy (acero1984):

I'll try

OpenStudy (anonymous):

After you find the two numbers, find the difference between them, and there's your answer. I didn't finish explaining before.

OpenStudy (anonymous):

@beckyg111